OpenStudy (anonymous):

Solve: -1/2x+3=-x+7 Oh by the way -1/2 is a fraction.

OpenStudy (anonymous):

\[\frac{ -1 }{ 2 }x + 3 = -x + 7\] just rearrange, get the like terms together and solve!

OpenStudy (anonymous):

\[\frac{ -1 }{ 2 }x + x = 7 - 3\] and solve for x

OpenStudy (anonymous):

why exactly would I rearrange if you don't mind me asking?

OpenStudy (primeralph):

you have mixed types of values on either side You can't add 2 apples and 2 oranges; only apples and oranges or oranges and oranges

OpenStudy (anonymous):

x=8